OpenStudy (anonymous):

x divided by - 4 - (-8) = 12. 16 80 -80 -16

OpenStudy (anonymous):

could you be a little more clear as to what the problem is? is this right: (x/-4)-(-8)=12?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

cn you rewite the equation?

myininaya (myininaya):

\[\frac{x}{-4}-(-8)=12\] option 1 \[\frac{x}{-4-(-8)}=12\] option 2 which option is your problem?

myininaya (myininaya):

you can also say none of the above

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

\[\frac{-x}{4}-(-8)=12\] x= (12-8)/*(-4) x =4*-4 x =-16

OpenStudy (anonymous):

x= (12-8)*(-4) x =4*-4 x =-16

OpenStudy (anonymous):

click good answer if I desever

jhonyy9 (jhonyy9):

x=-16